# Model Lock [Uncaptioned image]: Locking Your Model With a Spell  [Uncaptioned image]

### v2

@@ 4.4 Ablation Studies
Table 7: A list of surrogate prompts (generated by GPT-4V based on the leaked edited images) and their unlocking performance ($\text{UP}_{sp}(\%)$). The experiments were conducted with MAE-pretrained ViT-base and the Pets dataset.
Victim Model Surrogate Prompts $\text{UP}_{sp}(\%)$
Increase the sharpness and clarity of a petś facial features. 63.21
Style lock w Lora Dreamlike texture overlay on pet photos. 30.25
Enhanced saturation with a surreal touch for animal portraits. 7.93
Oil painting effect applied to dogs and cats photography. 7.69
Abstract color infusion for a whimsical look in pet images. 61.30
4.5 Robustness to Adaptive Attacks
Here, we test the robustness of ModelLock to adaptive attacks. We assume the attacker knows the open-source diffusion model and its detailed setup for ModelLock but not the key prompt. We also assume the defender can download different LoRA-tuned weights from Hugging Face or Civitai to make the editing model more covert. That is, the attacker only knows the diffusion model but not the LoRA module. Following this assumption, we test two adaptive attack scenarios: (1) the edited images are not leaked, and (2) a few edited images are leaked. In case (1), the attacker cannot obtain the edited data, so it has to guess the key prompt. We call this type of attack Random Prompt (RP). In case (2), the attacker can access part of the edited data and use the data to estimate the key prompt with the help of GPT-4V. We call this type of attack Surrogate Prompt (SP). We are interested in the unlocked performance by the two attacks which are denoted as $\text{UP}_{rp}$ and $\text{UP}_{sp}$, respectively. Note that the use of LoRA by the defender may slightly decrease the model’s task performance.
Figure 6: Unlocked accuracy ($\text{UP}_{rp}$, %) by 5 types of random prompts on Pets dataset. We test models locked using pre-trained Stable Diffusion with or without (w/wo) Lora.
@@ 5 Conclusion
While showing promising results in different vision tasks, our ModelLock is only a preliminary exploration of the model lock paradigm. There is still much room for improvement in terms of choosing the right key prompt or using optimization strategies to reduce the negative impact on the unlocked performance. We will also explore more fine-grained locking strategies that can release varying performance levels given different keys.
@@
Table 8: Candidate prompts from GPT4 for testing different kinds of Style lock ModelLock’s Unlocked Performance (UP) and Locked Performance (LP).
When generating Irrelevant Object and Relevant Object with GPT4-V, we provide a list of labels of ImageNet-1K as a reference. In Table 9, the parentheses following each returned candidate object prompt indicate the corresponding ImageNet-1K label index. Interestingly, we find that GPT4-V may produce object prompts that are not in ImageNet-1K (e.g., seahorse), suggesting that the range of potential key prompts is quite extensive.
@@
Table 9: Candidate prompts from GPT4 for testing different relevant and irrelevant Object lock ModelLock’s Unlocked Performance (UP) and Locked Performance (LP).
@@ Appendix B Example edited images with different prompts
Figure 7: Examples of edited images by Style, Object, and Combination of symbols lock prompts. All examples shown here are edited images synthesized with blending ratio $\gamma=0.5$ on the raw images after the diffusion editing process.
As can be observed in Figure 7, the edited images exhibit certain discrepancies compared to the raw images, although these differences are not pronounced. Particularly for edited images using prompts such as ”adding irrelevant object in the background (e.g., Basketball)” and ”using a sequence of characters (e.g., ”#9T8oN07y1%6\$”)”, the alterations are nearly imperceptible. One reason for this is that we employed significantly fewer inference steps (e.g., inference step = 5) than typical image editing. This has two advantages: 1) it improves the stealthiness of ModelLock, and 2) it reduces the computational cost of image editing.
@@
Appendix C Example edited images with traditional backdoor patterns
As shown in Figure 8, the patch pattern of BadNets and the predetermined pattern ”Hello Kitty” of Blend cause the edited samples to be distinguishable from the raw images. In the adaptive attack scenarios described in Section 4.5, when a few edited images are leaked, these patterns can be easily perceived and exploited by the attacker to unlock the model, undermining the security of ModelLock. However, the modifications by our ModelLock are minimized by the diffusion model, rendering them virtually imperceptible. Arguably, without the raw images, the edited images by ModelLock do not stand out conspicuously.
Figure 8: Example edited images using traditional backdoor patterns (BadNets and Blend) as locks. Note that BadNets-5, BadNets-20, and BadNets-50 correspond to patch trigger sizes of 5, 20, and 50, respectively.

### v3

@@ 4.4. Ablation Studies
This indicates that the locking effect of the model deteriorates rapidly when the edited proportion exceeds a certain threshold (not much data left in the locking subset $\mathcal{D}_{L}$. Therefore, a trade-off is necessary when increasing the proportion of the edited dataset between the benefits brought by UP and the damage to the LP of the model.
4.5. Robustness to Adaptive Attacks
Here, we test the robustness of ModelLock to adaptive attacks. We assume the attacker knows the open-source diffusion model and its detailed setup for ModelLock but not the key prompt. We also assume the defender can download different LoRA-tuned weights from Hugging Face or Civitai to make the editing model more covert. That is, the attacker only knows the diffusion model but not the LoRA module.
Following the above assumption, we test two adaptive attack scenarios: (1) the edited images are not leaked, and (2) a few edited images are leaked. In case (1), the attacker cannot obtain the edited data, so it has to guess the key prompt. We call this attack Random Prompt (RP). In case (2), the attacker can access part of the edited data and use the data to estimate the key prompt with the help of GPT-4V. We call this attack Surrogate Prompt (SP). We are interested in the unlocked performance by the two attacks which are denoted as $\text{UP}_{rp}$ and $\text{UP}_{sp}$, respectively. Note that the use of LoRA by the defender may slightly decrease the model’s task performance.
@@ 5. Discussion and Conclusion
While showing promising results in different vision tasks, our ModelLock is only a preliminary exploration of the model lock paradigm. The remaining challenges include how to select or optimize the key prompt and how to apply it to tasks that are sensitive to input editing, such as detecting modification of real images. Promising future research directions include 1) Quantifying the distribution shift and subject semantic similarity to establish a threshold for candidate key prompts selection based on the editing results; 2) Using optimized editing with task-related constraints, such as employing an originality-preserving constraint to maintain the original properties of the image, or even integrating the lock design and editing model optimization processes into the objective function, such as using the loss of clean data with unperturbed raw labels for gradient ascent to form a bilevel optimization framework. It is also interesting to explore more fine-grained locking strategies that can release varying performance levels given different keys.
Figure 6. Unlocked accuracy ($\text{UP}_{rp}$(%)) by 5 types of random prompts on Pets dataset. We test models locked using pre-trained Stable Diffusion with or without (w/wo) LoRA.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant No. 62276067.
@@ Acknowledgements.
Table S1. Candidate prompts from GPT4-V for testing different types of Style locks in terms of the Unlocked Performance (UP) and Locked Performance (LP) of ModelLock.
@@
When generating Irrelevant Object and Relevant Object with GPT4-V, we provide a list of labels of ImageNet-1K as a reference. In Table S2, the parentheses following each returned candidate object prompt indicate the corresponding ImageNet-1K label index. Interestingly, we find that GPT4-V may produce object prompts that are not in ImageNet-1K (e.g., seahorse), suggesting that the range of potential key prompts is quite extensive.
Table S2. Candidate prompts from GPT4-V for testing different types of relevant and irrelevant Object locks in terms of the Unlocked Performance (UP) and Locked Performance (LP) of ModelLock.
@@
Table S3. Performances (UP/LP accuracies) of the Modellock locked model (using combinations of chrachters devoid of semantic content as key prompt) on Pets dataset. R50 stands for Resnet50.
Pre-trained Meaningless Prompt Pets
[…]
MAE with ’@de%^fwd#oi’ in the background 91.55/52.74
#9T8oN07y1%6\$ 91.66/58.60
SimCLR-R50 with ’@de%^fwd#oi’ in the background 84.63/32.03
#9T8oN07y1%6\$ 84.03/25.16
Table S4. A list of surrogate prompts (generated by GPT-4V based on the leaked edited images) and their unlocking performance ($\text{UP}_{sp}(\%)$). The experiments were conducted with MAE-pretrained ViT-base and the Pets dataset.
Victim Model Surrogate Prompts $\text{UP}_{sp}(\%)$
Increase the sharpness and clarity of a petś facial features. 63.21
Style lock w Lora Dreamlike texture overlay on pet photos. 30.25
Enhanced saturation with a surreal touch for animal portraits. 7.93
Oil painting effect applied to dogs and cats photography. 7.69
Abstract color infusion for a whimsical look in pet images. 61.30
@@ Appendix B Example edited images with different prompts
Figure S1. Examples of edited images by Style, Object, and Combination of symbols lock prompts. All examples shown here are edited images synthesized with blending ratio $\gamma=0.5$ on the raw images after the diffusion editing process.
As can be observed in Figure S1, the edited images exhibit certain discrepancies compared to the raw images, although these differences are not pronounced. Particularly for edited images using prompts such as ”adding irrelevant object in the background (e.g., Basketball)” and ”using a sequence of characters (e.g., ”#9T8oN07y1%6\$”)”, the alterations are nearly imperceptible. One reason for this is that we employed significantly fewer inference steps (e.g., inference step = 5) than typical image editing. This has two advantages: 1) it improves the stealthiness of ModelLock, and 2) it reduces the computational cost of image editing.
@@
Appendix C Experimental setup and example images of the traditional backdoor lock
We use ‘BadNets-5’, ‘BadNets-20’, and ‘BadNets-50’ to denote trigger patch sizes of 5, 20, and 50, respectively. For the Blend attack, the blending ratio is set to 8:2 (image:pattern) to ensure a more effective locking effect. For the WaNet attack, we slightly modified its setting into a data poison-only backdoor setting (following prior work (Huang et al., 2023)) to guarantee the uniformity of the model training process. The grid size for WaNet is set to $k=32$.
As shown in Figure S2, the patch pattern of BadNets and the predetermined pattern ”Hello Kitty” of Blend cause the edited samples to be distinguishable from the raw images. In the adaptive attack scenarios described in Section 4.5, when a few edited images are leaked, these patterns can be easily perceived and exploited by the attacker to unlock the model, undermining the security of ModelLock. However, the modifications by our ModelLock are minimized by the diffusion model, rendering them virtually imperceptible. Arguably, without the raw images, the edited images by ModelLock do not stand out conspicuously.
Figure S2. Example edited images using traditional backdoor patterns (BadNets and Blend) as locks. Note that BadNets-5, BadNets-20, and BadNets-50 correspond to patch trigger sizes of 5, 20, and 50, respectively.
Appendix D Several questions about Modellock
Q1. What are the differences between Modellock and alternative model protecting approaches?
Model lock is necessary when fine-tuning private models on Machine Learning as a Service (MLaaS) platforms to protect the model and training data from dishonest service providers. The private model/data owner can utilize a customized LoRA-equipped diffusion model to pre-process the data before uploading it to the MLaaS platform. And the trained (locked) model is unusable even if the service provider leaks it to the model/data owner’s competitors. It is also important to note that model encryption and model lock are two distinct types of model protection methods, and our ModelLock is compatible with model encryption.
Q2. Can the key be stolen?
It is difficult to reverse-engineer the key without the (clean - edited) image pairs due to the large search space of text tokens. Additionally, even with the correct key prompt, the model cannot be unlocked without knowing the exact diffusion model used (equipped with customized LoRA weights). Nonetheless, more strong key prompt attacks such as Monte Carlo optimization or repeated attack sequences being optimized should be further explored on our ModelLock.
Q3. Can Modellock be applied to LLM and MLM?
Despite LLMs’ superiority in general domains, large models perform only mediocrely in specialized and private domains (e.g., medical, financial, self-driving, embodied AI) due to a lack of training data. Therefore, for LLMs, ModelLock aims to capitalize on the lack of generalizability in these specialized domains to lock proprietary LLMs. Therefore, Modellock is still effective for LLM and MLM.
Figure S3. Left: Unlocked accuracy (UP) and locked accuracy (LP) of models fine-tuned on the Pets dataset using 70 different key prompts of 6 different types. Right: Pixel value distributions of the raw (purple) and edited (pink vs. green) datasets, with green editing are better model locks.